OpenStudy (anonymous):

f(x)=x^(-.5) how do I rewrite this with a positive radical?

OpenStudy (anonymous):

x^-a = 1/x^a

OpenStudy (anonymous):

f(x) = 1/sqrt(x)

OpenStudy (anonymous):

Can you see what it will be from what i wrote?

OpenStudy (anonymous):

is it \[1/x^{.5}\]then?

OpenStudy (anonymous):

.5 = 1/2 right so x^.5= x^1/2 and we know that x^a/b =\[\sqrt[a]{x^{b}}\]

OpenStudy (anonymous):

so its going to be 1/ squareroot (x)

OpenStudy (anonymous):

Algebraically how do I get rid of a sqrt that is on the bottom of a fraction? I haven't taken any math for like 2 years so I cant remember some stuff haha. for example am doing the limit at a point so I will have  how do I get rid of the square root to solve for h?